Please Help:
Find the future value of the ordinary annuity. Interest is compounded annually, unless otherwise indicated.
R = $900, i = 7% interest compounded semiannually for 10 years.
A) $23,637.33
B) $23,721.46
C) $25,451.71
D) $51,166.00

Answer by stanbon(75887)

R = $900, i = 7% interest compounded semiannually for 10 years.
---
A(t) = P(1 + (r/n))^(nt)
A(10) = 900(1 + (0.07/2))^(2*10)
A(10) = 900(1.035)^(20)
A(10) = 900*1.9897888...
A(10) = $1790.81
---------------
Comment:
You may have a typo in the posted problem.
-----------------
Cheers,
Stan H.

A) $23,637.33
B) $23,721.46
C) $25,451.71
D) $51,166.00